OpenStudy (anonymous):

if (x:41) ----> remainder=41, if ((x^-1) : 82) ---> remainder=81 if ((x^4-1) : 164) ---> remainder=163 The minimal value of x is ....?

OpenStudy (anonymous):

if ((x^2-1) : 82) ---> remainder=81 ?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

oh sorry, the first if (x:41) ----> remainder=40

OpenStudy (anonymous):

i think its impossible if (x:41) ----> remainder=40 then x=41k+40 k=0,1,2,... we have 2 cases case I) k=2m+1 m=0,1,2,... then x=41(2m+1)+40=82m+81 so (x^2-1=(x-1)(x+1)=(82m+80)(82m+82):82) ----> remainder=0 case II) k=2m m=0,1,2,... then x=41(2m)+40=82m+40 so (x^2-1=(x-1)(x+1)=(82m+39)(82m+41):82) ----> remainder=39*41 or 41

OpenStudy (anonymous):

let me know what is your idea